OpenStudy (anonymous):

Let S be a infinite set and let T be a finite subset of S. Prove that S − T is infinite.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

i dont know

OpenStudy (anonymous):

@mukushla

OpenStudy (dmnts):

\[(S)=(x _{1},x _{2},x _{3},...,x _{n-1},x _{n},x _{n+1},...)\] \[(T _{n})=(x _{1},x _{2},x _{3},...,x _{n-1},x _{n}), T _{n}\subset S\]then\[(S)-(T _{n})=(x _{1},x _{2},x _{3},...,x _{n-1},x _{n},x _{n+1},...)-(x _{1},x _{2},x _{3},...,x _{n-1},x _{n})\] \[(S)-(T _{n})=(x _{n+1},x _{n+2},...)\]

OpenStudy (anonymous):

I understand, thank you.

OpenStudy (anonymous):

I think sometimes in higher level mathematics we tend to over think things way too much.

OpenStudy (john_es):

I was thinking in an alternative proof in the line tukajo was probably thinking at first. We assume \[T\cup (S-T)=S, \ S \cap T=\phi \] Then we could think in the number of elements, or cardinal, \[Card (T)+Card(S-T)=Card(S)\] As T is finite, for some M, Card(T)<M. As S is infinite, for some N, Card(S)>N. So, suppose S-T is finite, then Card(S-T)<P, for some P. Then, \[Card(T)+Card(S-T)<M+P\Rightarrow Card(S)<M+P\] But this is in contradiction with our assumption, Card(S)>N, for some N.

OpenStudy (anonymous):

Yes John, that is a bit more geared towards what I had thought. Thank you, do you have time in a moment to assist in developing my thoughts on one more problem?